# The pattern block method for generating random numbers : Reformulation and generalization of the Ziggurat method using conditional random variables

### 19 Mar 2026 v2

@@ Abstract
The Ziggurat method is an efficient rejection sampling method for generating 11-dimensional normally distributed random numbers. This paper proposes the pattern block method, which is a generalization of the Ziggurat method. The pattern block method can be used to generate random numbers from multimodal density functions or multidimensional distributions. We demonstrate the effectiveness of the pattern block method through several examples.
Keywords: Random number generation, The Ziggurat method, The pattern block method. 0002020 Mathematics Subject Classification: Primary 65C10; Secondary 62-08.
@@ 1 Introduction
Random number generation is a foundational component of Monte Carlo methods. Recently, [1, 2, 3, 4, 5, 6] developed the Ziggurat method for generating 11-dimensional normally distributed random numbers and demonstrated its effectiveness through numerical examples. However, the random variable underlying the Ziggurat algorithm has not been defined. Therefore, in this study, we reformulate and generalize the Ziggurat method using conditional random variables. Conditional random variables have played an important role in [7, 8, 9, 10].
This paper is organized as follows. In Section 2, we present the pattern block method proposed in this study. Sections 3 and 4 present applications of the pattern block method. Section 5 concludes.
@@ 2 The pattern block method
Let $(E,\mathcal{E},\mu)$ be a $\sigma$-finite measure space. Let $f:E\rightarrow\mathbb{R}$ be an $\mathcal{E}/\mathcal{B}(\mathbb{R})$-measurable function that satisfies
@@
where $\mathcal{B}(\mathbb{R})$ is the Borel $\sigma$-algebra on $\mathbb{R}$. Then $\frac{1}{K}f(x)$ is a probability density function. Let
@@
Let $\widetilde{\pi}:E\times\mathbb{R}\to E$ be the projection defined by
@@
Let $N\in\mathbb{N}$. Assume that $B_{1},B_{2},\cdots,B_{N}\in\mathcal{E}\otimes\mathcal{B}(\mathbb{R})$ and $B:=\bigcup_{i=1}^{N}B_{i}$ satisfy the following three conditions:
@@
In this study, we call $B_{1},B_{2},\cdots,B_{N}$ “pattern block of f(x)f(x).” Using (1), (4), (5), we obtain
@@
Figure 1: $f(x)$ and its pattern block
Let $\Psi$ be the block selection function defined by
@@
Let $\Lambda$ be the acceptance set defined by
@@
Let $\pi_{\Psi}:[0,1)\times E^{N}\times\mathbb{R}^{N}\to E$ be the projection
@@
Let $Z:(\Omega,\mathcal{F},P)\rightarrow(\mathbb{R}\times E^{N}\times\mathbb{R}^{N},\ \mathcal{B}(\mathbb{R})\otimes\mathcal{E}\otimes\mathcal{E}\otimes\cdots\otimes\mathcal{E}\otimes\mathcal{B}(\mathbb{R}^{N}))$ be the random variable given by
@@
Using (5), (14) and the Fubini’s theorem, we obtain (12). ∎
By Lemma 2.1, (1) and (6), it holds that
@@
Then $Z|_{\Lambda}$ is a random variable. Let $X:(Z^{-1}(\Lambda),Z^{-1}(\Lambda)\cap\mathcal{F},P_{Z^{-1}(\Lambda)})\to(E,\mathcal{E})$ be the random variable defined by
@@
$Z|_{\Lambda}$ and $X$ are random variables defined on $(Z^{-1}(\Lambda),Z^{-1}(\Lambda)\cap\mathcal{F},P_{Z^{-1}(\Lambda)})$, and are called conditional random variables. Using Lemma 2.1 and (15), we obtain the following theorem.
@@
Thus, using the random variable $\pi_{\Psi}(Z|_{\Lambda})$, we can generate random numbers from the probability density function $\frac{1}{K}f(x)$. We call this technique the pattern block method. The algorithm of the pattern block method is as follows:
@@
Let $R_{{\rm adoption}}$ be the adoption rate defined by
@@ 3 Applications to generate 11-dimensional random numbers
In this section, we assume that $(E,\mathcal{E})$ is $(\mathbb{R},\mathcal{B}(\mathbb{R}))$ and $\mu$ is the 11-dimensional Lebesgue measure on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$. Then $\nu$ defined by (2) is the 22-dimensional Lebesgue measure on $(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2}))$. For every $C\in\mathcal{B}(\mathbb{R}^{2})$, $|C|$ denotes the 22-dimensional Lebesgue measure $\nu(C)$.
@@ 3.1 The Ziggurat method
In this subsection, we explain that the pattern block method is a generalization of the Ziggurat method. Assume that a probability density function $f(x)$ is strictly decreasing for $x>0$ and satisfies $\int_{0}^{\infty}f(x)\ dx=1$. Let $N\in\mathbb{N}$ with $N\geq 2$, and let $0=x_{0}<x_{1}<\cdots<x_{N-1}$. Let
@@
be the pattern block of $f(x)$. Then the pattern block method corresponds to the Ziggurat method.
3.2 The density function with singularities
Let $\varphi(x)=\dfrac{1}{\pi\sqrt{x(1-x)}}$ and $h(x)=1+\sin(8\pi x)$. Let $f(x)$ be the density function
@@ 3.2 The density function with singularities
Let $N=8$. We define the pattern block of $f(x)$ as
@@
Figure 3: $f(x)$ and its pattern block $B_{1},B_{2},\cdots,B_{8}$
Since $\varphi(x)$ is the density function of the Beta distribution $B\left(\tfrac{1}{2},\tfrac{1}{2}\right)$, its cumulative distribution function $\Phi(x)$ is given by
@@
Assume that the independent random variables $U$, $\xi$, $\eta$ defined on a probability space $(\Omega,\mathcal{F},P)$ are uniformly distributed on $[0,1)$. We define the random variables $(V_{i},W_{i})$ $(i=1,2,\cdots,8)$ as
@@
Then $(V_{i},W_{i})$ $(i=1,2,\cdots,8)$ satisfy (9). The adoption rate can be calculated as follows:
@@
Thus, by applying the pattern block method, we obtain $10,000$ random numbers from the probability density function $f(x)$. To obtain $10,000$ random numbers, $15,048$ random number triplets $(U,\xi,\eta)$ were required. Fig. 4 shows the density scale histogram of $10,000$ random numbers.
@@ 4 Application to generate 22-dimensional random numbers
In this section, we assume that $(E,\mathcal{E})$ is given by $E=[\underline{a},\overline{a}]\times[\underline{a},\overline{a}]$ and $\mathcal{E}=\mathcal{B}(E)$, where $\underline{a}=-4$ and $\overline{a}=4$. Let $\mu$ be the 22-dimensional Lebesgue measure on $(E,\mathcal{E})$. Then $\nu$ is the 33-dimensional Lebesgue measure on $(E\times\mathbb{R},\mathcal{B}(E\times\mathbb{R}))$. For every $C\in\mathcal{B}(E\times\mathbb{R})$, $|C|$ denotes the 33-dimensional Lebesgue measure $\nu(C)$. Let $f(x_{1},x_{2})$ be the density function on $(E,\mathcal{E})$
@@
Let $N=5$. We define the pattern block of $f(x_{1},x_{2})$ as $B_{i}=E_{i}\times\left[\underline{\kappa}_{i},\overline{\kappa}_{i}\right]$ $(1\leq i\leq 5)$, where
@@
Assume that the independent random variables $U$, $\xi_{1}$, $\xi_{2}$, $\eta$ defined on a probability space $(\Omega,\mathcal{F},P)$ are uniformly distributed on $[0,1)$. Let
@@
Then $(V_{i}^{(1)},V_{i}^{(2)},W_{i})$ $(i=3,4,5)$ satisfy (9). Thus, by applying the pattern block method, we obtain $500,000$ random vectors from the probability density function $f(x_{1},x_{2})$. To obtain $500,000$ random vectors, $1,375,334$ random number quartets $(U,V_{\Psi(U)}^{(1)},V_{\Psi(U)}^{(2)},W_{\Psi(U)})$ were required. The adoption rate can be calculated as follows:
@@
Fig. 9 and Fig. 9 show the density scale histogram of $500,000$ random vectors and the density function $y=f(x_{1},x_{2})$, respectively.
@@ 5 Conclusion
We introduced the pattern block method, which is a generalization of the Ziggurat method. In our future work, we aim to develop an algorithm that identifies pattern block of a given density function with high adoption rate and low search cost.
@@ Acknowledgments
The authors would like to thank Prof. Masaaki Fukasawa (Osaka University) and Prof. Suguru Yamanaka (Aoyama Gakuin University) for their helpful comments and discussions on the subject matter. This work was supported by JSPS KAKENHI Grant Numbers JP22K01556 and JP25K05168.

### 24 Mar 2026 v3

# Pattern block method for generating random numbers : Reformulation and generalization of the Ziggurat method using conditional random variables

@@ Abstract
The Ziggurat method is an efficient rejection sampling technique for generating one-dimensional normally distributed random numbers. This study proposes the pattern block method, a generalization of the Ziggurat method. The pattern block method enables the generation of random numbers from multimodal density functions and multidimensional distributions. The effectiveness of the pattern block method is demonstrated through several examples.
Keywords: Random number generation, Ziggurat method, Pattern block method. 0002020 Mathematics Subject Classification: Primary 65C10; Secondary 62-08.
@@ 1 Introduction
Random number generation is a fundamental component of Monte Carlo methods. Recently, [1, 2, 3, 4, 5, 6] developed the Ziggurat method for generating 11-dimensional normally distributed random numbers and demonstrated its effectiveness using numerical examples. However, the random variable underlying the Ziggurat algorithm has not yet been formally defined. Therefore, this study reformulates and generalizes the Ziggurat method using conditional random variables. Conditional random variables play an important role in [7, 8, 9, 10].
The remainder of this paper is organized as follows. Section 2 presents the pattern block method proposed in this study. Sections 3 and 4 present applications of the pattern block method. Finally, Section 5 concludes.
@@ 2 The pattern block method
Let $(E,\mathcal{E},\mu)$ be a $\sigma$-finite measure space. Let $f:E\rightarrow\mathbb{R}$ be a $\mathcal{E}/\mathcal{B}(\mathbb{R})$-measurable function that satisfies
@@
where $\mathcal{B}(\mathbb{R})$ denotes the Borel $\sigma$-algebra of $\mathbb{R}$. Then $\frac{1}{K}f(x)$ is a probability density function. Let
@@
Let $\widetilde{\pi}:E\times\mathbb{R}\to E$ be a projection defined by
@@
Let $N\in\mathbb{N}$. We assume that $B_{1},B_{2},\cdots,B_{N}\in\mathcal{E}\otimes\mathcal{B}(\mathbb{R})$ and $B:=\bigcup_{i=1}^{N}B_{i}$ satisfy the following three conditions:
@@
We refer to $B_{1},B_{2},\cdots,B_{N}$ as the pattern blocks of $f(x)$. Using (1), (4), and (5), we obtain
@@
Figure 1: $f(x)$ and its pattern blocks
Let $\Psi$ be the block selection function defined as
@@
Let $\Lambda$ be the acceptance set defined as
@@
Let $\pi_{\Psi}:[0,1)\times E^{N}\times\mathbb{R}^{N}\to E$ denote the projection:
@@
Let $Z:(\Omega,\mathcal{F},P)\rightarrow(\mathbb{R}\times E^{N}\times\mathbb{R}^{N},\ \mathcal{B}(\mathbb{R})\otimes\mathcal{E}^{\otimes N}\otimes\mathcal{B}(\mathbb{R}^{N}))$ be a random variable given by
@@
Using (5), (14), and Fubini’s theorem, we obtain (12). ∎
By Lemma 2.1, (1), and (6), it follows that
@@
Thus, $Z|_{\Lambda}$ is a random variable. Let $X:(Z^{-1}(\Lambda),Z^{-1}(\Lambda)\cap\mathcal{F},P_{Z^{-1}(\Lambda)})\to(E,\mathcal{E})$ be a random variable defined as follows:
@@
$Z|_{\Lambda}$ and $X$ are conditional random variables defined on $(Z^{-1}(\Lambda),Z^{-1}(\Lambda)\cap\mathcal{F},P_{Z^{-1}(\Lambda)})$. Using Lemma 2.1 and (15), we obtain the following theorem:
@@
Thus, using the conditional random variable $\pi_{\Psi}(Z|_{\Lambda})$, we can generate random numbers from the density function $\frac{1}{K}f(x)$. We call this technique the pattern block method. The algorithm is as follows:
@@
Let $R_{{\rm adoption}}$ be the adoption rate defined as
@@ 3 Applications to generate 11-dimensional random numbers
In this section, we assume the following. $(E,\mathcal{E})$ is $(\mathbb{R},\mathcal{B}(\mathbb{R}))$, and $\mu$ is the 11-dimensional Lebesgue measure on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$. Then, $\nu$, defined by (2), is the 22-dimensional Lebesgue measure on $(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2}))$. For every $C\in\mathcal{B}(\mathbb{R}^{2})$, $|C|$ denotes the 22-dimensional Lebesgue measure $\nu(C)$.
@@ 3.1 The Ziggurat method
In this subsection, we explain that the pattern block method is a generalization of the Ziggurat method. We assume that the probability density function $f(x)$ strictly decreases for $x>0$ and satisfies $\int_{0}^{\infty}f(x)\ dx=1$. Let $N\in\mathbb{N}$ with $N\geq 2$, and let $0=x_{0}<x_{1}<\cdots<x_{N-1}$. Let
@@
be the pattern blocks of $f(x)$. The pattern block method then corresponds to the Ziggurat method.
3.2 Density function with singularities
Let $\varphi(x)=\dfrac{1}{\pi\sqrt{x(1-x)}}$ and $h(x)=1+\sin(8\pi x)$. Let $f(x)$ be the density function defined as
@@ 3.2 Density function with singularities
Let $N=8$. The pattern blocks of $f(x)$ are defined as
@@
Figure 3: $f(x)$ and its pattern blocks $B_{1},B_{2},\cdots,B_{8}$
Because $\varphi(x)$ is the density function of the Beta distribution $B\left(\tfrac{1}{2},\tfrac{1}{2}\right)$, its cumulative distribution function $\Phi(x)$ is given by
@@
We assume that the independent random variables $U$, $\xi$, and $\eta$, defined on the probability space $(\Omega,\mathcal{F},P)$, are uniformly distributed on $[0,1)$. We define the random variables $(V_{i},W_{i})$ $(i=1,2,\cdots,8)$ as
@@
Subsequently, $(V_{i},W_{i})$ $(i=1,2,\cdots,8)$ satisfy (9). The adoption rate is calculated as follows:
@@
Thus, by applying the pattern block method, we obtained $10000$ random numbers from the probability density function $f(x)$. To obtain $10000$ random numbers, $15048$ random number triplets $(U,\xi,\eta)$ were required. Fig. 4 shows the density scale histogram for $10000$ random numbers.
@@ 4 Application to generate 22-dimensional random numbers
In this section, we assume that $(E,\mathcal{E})$ is defined as follows: $E=[\underline{a},\overline{a}]\times[\underline{a},\overline{a}]$ and $\mathcal{E}=\mathcal{B}(E)$, where $\underline{a}=-4$ and $\overline{a}=4$. Let $\mu$ be the 22-dimensional Lebesgue measure on $(E,\mathcal{E})$. Then $\nu$ is the 33-dimensional Lebesgue measure on $(E\times\mathbb{R},\mathcal{B}(E\times\mathbb{R}))$. For every $C\in\mathcal{B}(E\times\mathbb{R})$, $|C|$ denotes the 33-dimensional Lebesgue measure $\nu(C)$. Let $f(x_{1},x_{2})$ be the density function on $(E,\mathcal{E})$:
@@
Let $N=5$. The pattern blocks of $f(x_{1},x_{2})$ are defined as $B_{i}=E_{i}\times\left[\underline{\kappa}_{i},\overline{\kappa}_{i}\right]$ $(1\leq i\leq 5)$, where
@@
We assume that the independent random variables $U$, $\xi_{1}$, $\xi_{2}$, and $\eta$, defined on the probability space $(\Omega,\mathcal{F},P)$, are uniformly distributed on $[0,1)$. Let
@@
Subsequently, $(V_{i}^{(1)},V_{i}^{(2)},W_{i})$ $(i=3,4,5)$ satisfy (9). Thus, by applying the pattern block method, we obtained $500000$ random vectors from the probability density function $f(x_{1},x_{2})$. To obtain $500000$ random vectors, $1375334$ random number quartets $(U,V_{\Psi(U)}^{(1)},V_{\Psi(U)}^{(2)},W_{\Psi(U)})$ were required. The adoption rate is calculated as follows:
@@
Figs. 9 and 9 show the density scale histogram for $500000$ random vectors and the density function $y=f(x_{1},x_{2})$.
@@ 5 Conclusion
This study introduced the pattern block method, a generalization of the Ziggurat method. Future work will focus on developing an algorithm that identifies the pattern blocks of a given density function with high adoption rate and low search cost.
@@ Acknowledgments
The authors thank Prof. Matsuura Yusuke (National Institute of Advanced Industrial Science and Technology), Prof. Masaaki Fukasawa (Osaka University) and Prof. Suguru Yamanaka (Aoyama Gakuin University) for their helpful comments and discussions on this study. We also thank Editage (www.editage.com) for English language editing. This work was supported by JSPS KAKENHI Grant Numbers JP22K01556 and JP25K05168.